Margy is trying to improve her cardio endurance by performing an exercise in which she alternates walking and running 100.0 m each. If Margy is walking at 1.4 m/s and accelerates at 0.20 m/s2 during one of the running portions, what is her final velocity at the end of the 100.0 m? Round your answer to the nearest tenth.
1
Expert's answer
2020-08-19T12:57:04-0400

"v_f^2-v_i^2=2as\\v_f^2-1.4^2=2(0.2)(100)\\v_f=6.5\frac{m}{s}"